# 360VFI: A Dataset and Benchmark for Omnidirectional Video Frame Interpolation

### v2

@@ Abstract
With the development of VR-related techniques, viewers can enjoy a realistic and immersive experience through a head-mounted display, while omnidirectional video with a low frame rate can lead to user dizziness. However, the prevailing plane frame interpolation methodologies are unsuitable for Omnidirectional Video Interpolation, chiefly due to the lack of models tailored to such videos with strong distortion, compounded by the scarcity of valuable datasets for Omnidirectional Video Frame Interpolation. In this paper, we introduce the benchmark dataset, 360VFI, for Omnidirectional Video Frame Interpolation. We present a practical implementation that introduces a distortion prior from omnidirectional video into the network to modulate distortions. We especially propose a pyramid distortion-sensitive feature extractor that uses the unique characteristics of equirectangular projection (ERP) format as prior information. Moreover, we devise a decoder that uses an affine transformation to facilitate the synthesis of intermediate frames further. 360VFI is the first dataset and benchmark that explores the challenge of Omnidirectional Video Frame Interpolation. Through our benchmark analysis, we presented four different distortion conditions scenes in the proposed 360VFI dataset to evaluate the challenge triggered by distortion during interpolation. Besides, experimental results demonstrate that Omnidirectional Video Interpolation can be effectively improved by modeling for omnidirectional distortion.
@@ 1 Introduction
Figure 1: Left: Traditional Video Frame Interpolation: The inputs are two adjacent plane frames and the output is a target plane frame. Right: Omnidirectional Video Frame Interpolation: The inputs are two adjacent omnidirectional frames with a full field-of-view from an omnidirectional video and the output is a target omnidirectional frame. Original-formatted omnidirectional frames are spherical and the most common format is Equirectangular projection (ERP). The two kinds of omnidirectional video formats can be projected from each other and our proposed method tackled ERP videos.
Traditional Plane Video Frame Interpolation (Plane VFI) techniques have emerged as a basic solution to address the challenges of low frame rates. The development of optical flow networks (Dosovitskiy et al., 2015; Sun et al., 2018; Teed & Deng, 2020; Kong et al., 2021) has led to notable progress in flow-based Plane VFI methods (Jiang et al., 2018a; Xu et al., 2019; Niklaus & Liu, 2020; Park et al., 2021). This is attributed to the capability of optical flow to establish explicit correspondences for frame registration in video frame sequences. Flow-based methods typically follow a three-step process: First, estimate the optical flow between two input frames and an arbitrary target frame. Then warp the input frames or context features using predicted flow fields for spatial alignment. Finally, refine warped frames or features and generate the target frame using a synthesis network. Recently, IFRNet (Kong et al., 2022) was proposed to benefit intermediate flow and intermediate feature with each other, with the ability to generate sharper moving objects and capture better texture details. Similarly, in the omnidirectional domain, interpolating frames via optical flow is effective but not efficient. Omnidirectional optical flow estimation has been explored recently. In SLOF (Bhandari et al., 2022), ERP video frames were re-projected onto the sphere to apply rotational augmentation. In MPF (Li et al., 2022), the ERP format frames were transformed into multiple types of projection formats, and these different projection formats were fused to learn a more precise flow. However, these two works have not considered using ERP distortion as prior knowledge. Therefore, it is not efficient to use their estimated optical flow to interpolate omnidirectional frames. As shown in Figure 1, ODVs are typically stored and transmitted in equirectangular projection (ERP) format, which introduces distortions. So traditional frame interpolation methods struggle to handle effectively. Consequently, dealing with ERP video frames directly using distortion as prior knowledge is an effective and efficient way.
@@
Datasets for plane VFI tasks have been sufficiently researched, including Vimeo90K (Xue et al., 2019), UCF101 (Soomro et al., 2012), and SNU-FILM (Choi et al., 2020). Vimeo90K is the most popular dataset for plane VFI, which is composed of triplet samples. However, Vimeo90K is not designed for motion extent, and this is very important in Omnidirectional Video Interpolation because of its latitude-dependent distortion. SNU-FILM was proposed for the motion-based interpolation benchmark. Both are focused on Plane VFI. Recently, omnidirectional video datasets have caught the attention of the computer community. A few datasets were proposed for Omnidirectional Video Super-Resolution including ODV360 (Wang et al., 2023), 360VDS, and 360UHD (Baniya et al., 2023). They are collected from both Youtube and by themselves. Therefore, there are no Omnidirectional Video Frame Interpolation (Omni-VFI) datasets that have both sufficient and various samples.
@@
To alleviate these problems, we introduce a dataset for Omnidirectional Video Interpolation, called 360VFI dataset. Besides, we also propose a benchmark (360VFI), which injects the omnidirectional distortion into the network, based on the 360VFI dataset. Specifically, in the 360VFI dataset, our evaluation benchmark has four different settings. We classified the triplets based on vertical motion extent due to the latitude-dependent distortion in the omnidirectional frame. In 360VFI Network, we propose the DistortionGuard module to extract less distorted pyramid features from input frames. Moreover, we propose OmniFTB to gradually refine a distorted ERP target frame. Experimental results demonstrate that the Omnidirectional Video Interpolation can be effectively improved by modeling for omnidirectional distortion.
Overall, this paper makes the following contributions:
  • We present the first Omnidirectional Video Frame Interpolation (Omni-VFI) datasets 360VFI covering of various motion and content.
  • We propose a benchmark for evaluating Omni-VFI in different vertical motion, and our 360VFI Net successfully improves Omni-VFI performance.
@@ 2.1 Omnidirectional Data Processing
Omnidirectional images and videos have been attracting increasing attention from computer vision and graphics researchers. Due to the special representation of omnidirectional images, dedicated methods were developed for analysis and processing tasks, such as depth estimation (Wang et al., 2020a), salient object detection (Li et al., 2019b; Wu et al., 2022; Monroy et al., 2018; Zhang et al., 2018), edit propagation (Zhang et al., 2021), and video stabilization (Kopf, 2016; Tang et al., 2019). With the recent development of immersive technologies, researchers have been working on the specific problems of omnidirectional video-based applications, such as automatic 360° navigation (Kang & Cho, 2019; Hu et al., 2017), omnidirectional video assessment (Li et al., 2019a), and immersive video editing (Nguyen et al., 2017). However, the absence of consistent temporal correspondences poses challenges for various applications, including Omni-VFI. In response to this challenge, our method aims to address the lack of reliable temporal correspondences in omnidirectional video processing. Furthermore, significant research (Esteves et al., 2018; Cohen et al., 2018) has focused on omnidirectional or 360° data, including spherical and ERP data, aimed at learning sphere representations
@@ 2.2 Video Frame Interpolation
Video Frame Interpolation aims at generating non-existent frames between two adjacent input frames. In Long et al. (2016), a pioneering learning-based method was proposed that can directly synthesize intermediate frames between two frames. Consequently, frame interpolation methods based on spatially adaptive convolution kernels (Niklaus et al., 2017; Lee et al., 2020; Cheng & Chen, 2021) or pixel phases (Meyer et al., 2015; 2018) were proposed. However, the former leads to large parameters and high complexity, especially when dealing with complex motion. The latter has difficulty handling complex motion due to an inaccurate estimation of phase and amplitude values. After that, many flow-based methods (Jiang et al., 2018b; Xue et al., 2019) use optical flow to guide warping, but the inaccuracy of the predicted optical flow usually causes distortion of the result, so extra measures (Bao et al., 2019; 2018) are usually taken to refine the warped frame. Some methods generate intermediate frames directly like CAIN (Choi et al., 2020) using channel attention and RIFE (Huang et al., 2022) using distillation to get a good result from one simple architecture. IFRNet (Kong et al., 2022) proposes a novel single encoder-decoder-based method to jointly perform intermediate flow estimation and intermediate feature refinement for efficient VFI, which also performs real-time inference with excellent accuracy. All Video Frame Interpolation methods mentioned above cannot generate well on omnidirectional frames because they are not designed for them. Additionally, to our best knowledge, few methods have been proposed for Omni-VFI.
@@ 2.3 Conditional Modeling and Integration
In OSRT (Yu et al., 2023), a distortion-aware transformer used for omnidirectional image super-resolution was proposed. They take the ERP distortion priors as a condition in the attention mechanism to modulate distortion. Specifically, OSRT deforms the feature maps using offsets calculated from the latitude-dependent conditions, continuously and adaptively modulating and encoding the ERP distortion into the network, achieving modeling and correction of the ERP distortion. The Spatial Feature Transform (SFT) (Wang et al., 2018) uses semantic segmentation probability maps as categorical priors to provide useful conditioning information for the image super-resolution task. Specifically, the low-resolution input image is first passed through a semantic segmentation network to obtain the corresponding probability maps, which represent the probability of each pixel belonging to different semantic classes (e.g., sky, building, vegetation, etc.). Then, a novel Spatial Feature Transform (SFT) layer takes these probability maps as conditions, learns to generate scaling parameters specific to each semantic class, and uses them to modulate the intermediate feature maps. This allows the generated high-resolution image to have more realistic and intricate textural details consistent with the corresponding categories.
@@ 2.4 Omnidirectional Datasets
While datasets for plane VFI tasks have been extensively researched, including Vimeo90K (Xue et al., 2019), UCF101 (Soomro et al., 2012), and SNU-FILM (Choi et al., 2020), there is a lack of dedicated datasets for Omni-VFI. Leveraging existing datasets can streamline the benchmark creation process for Omni-VFI, ensuring standardization and reducing the learning curve for researchers. However, there is no dataset used in the computer vision community directly for Omni-VFI. Thanks to the contributions of Wang et al. (2023) and S3PO (Baniya et al., 2023), high-resolution ODV super-resolution datasets ODV360, 360VDS, and 360UHD have been introduced, which can be utilized in Omni-VFI.
@@ 3.2 Proposed 360VFI Dataset
Our 360VFI dataset adopts a similar format as Vimeo90K (Xue et al., 2019). Each sample in our dataset consists of a triplet of video frames. Notably, the first and third frames are designated as input frames for the Omni-VFI model, while the second serves as the ground truth target frame. In the ODV360 (Wang et al., 2023) dataset, each video comprises 100 frames, while in the 360VDS and 360UHD (Baniya et al., 2023) dataset, the videos consist of 20 frames. We omit the final frame from each video in the ODV360 dataset, resulting in 99 frames per video, which are subsequently divided into 33 triplets. For videos in the 360VDS and 360UHD datasets, the first and last frames are discarded, leaving 18 frames per video, which are divided into 6 triplets. Each triplet constitutes an independent sample and the whole dataset consists of 930 triplets. This comprehensive arrangement enables a holistic evaluation of interpolation algorithms against genuine ODV content. We randomly divide these videos into training and testing sets.
@@
Videos in our dataset cover various scenarios, including natural landscapes, playgrounds, interiors of houses and cars, and indoor markets, as shown in Figure 2. To facilitate nuanced evaluations and benchmarking, we stratified the dataset into four distinct settings based on the varying degrees of motion inherent within the omnidirectional scenes, ranging from 0.19 to 9.69. These settings are categorized as easy, middle, hard, and extreme; the triplet numbers and their respective storage are illustrated in Table 2. The motion ranges for these settings are as follows: [0.19, 2], [2, 3], [3, 4], and [4, 9.69]. The frames and optical flow of different partitions are illustrated in Figure 3. This stratification allows researchers to systematically assess the performance of their models across varying levels of motion complexity, facilitating a deeper understanding of model robustness and generalization capabilities. In Table $1$VFI: Video Frame Interpolation; VSR: Video Super-Resolution; OLSE: Outdoor Light Source Estimation; OD: Object Detection, we compare the 360VFI dataset with other datasets. These datasets include plane video, 360° video, and 360° images. Our dataset is the first one composed of 360° video for Omnidirectional Video Interpolation. Our dataset establishes a solid foundation for future research in Omni-VFI. By amalgamating diverse datasets and incorporating motion-based stratification, it not only enriches the existing repository of ODV data, but also sets a precedent for structured evaluation and benchmarking in this evolving field. We anticipate that our dataset will serve as a catalyst for innovation and foster the development of more sophisticated and resilient Omnidirectional Video Interpolation models in the future.
@@ 4 Method
where $\Psi$ defines the prior upon which the mapping function can condition. As mentioned in Section 2.3, there are many work using prior knowledge to guide their task in other computer vision tasks. Hence, we propose 360VFI Net, including DistortionGuard and OmniFTB.
@@
4.1 Distortions in Omnidirectional Video
ERP is a projection method used to map spherical surfaces onto a plane. In ERP, every point on the sphere is mapped onto the plane in such a way that each line from the polar center point maintains the same distance. So the distortion of ERP is latitude-dependent and all the pixels at the same height have the same distortion extent. This projection is frequently utilized in the creation of omnidirectional images and videos to depict environments more authentically and naturally, as it maintains depth and distance perception while minimizing distortion. As ODIs under different projection types are constrained by different transforming equations, the distortion caused by each type is inconsistent. ERP is the most convenient projection type for storage or transmission. To explain the causation of distortions more effectively, we follow the definition of stretching ratio ($\mathbf{K}$) in Sun et al. (2017a), which signifies the degree of distortion at various locations, measured from the desired projection type to the ideal spherical surface. $\mathbf{K}$ is determined by the variation in the area from one type of projection to another. As shown in Figure 5, the coordinate in ERP is defined as $x$ and $y$. ERP stretching ratio can be derived as
@@ 4.1 Distortions in Omnidirectional Video
where $\delta S$ represents the area on the spherical surface and $\delta P$ represents the area onthe projection plane, $x\in(-\pi,\pi)$, $y\in(-\frac{\pi}{2},\frac{\pi}{2})$.
From equation 5, we conclude that the ERP distortion is only determined by its degree of latitude. ERP distortion is the heaviest in the polar areas. ERP images suffer a distortion caused by a non-consistency area stretching ratio from an ideal spherical surface. Referred from equation 5, for a input $X\in\mathbb{R}^{C\times M\times N}$, the distortion condition map $C_{d}\in\mathbb{R}^{1\times M\times N}$ is derived by:
@@
where $m$ is the current height input frames.
Spatial distortion within image frames and the estimation of temporal motion patterns represent significant challenges in Omnidirectional Video Interpolation. In terms of spatial distortion, the distortion within image frames primarily arises from the projection methods used to capture omnidirectional scenes, such as Equirectangular Projection (ERP). Due to the unique characteristics of ERP, such as latitude-dependent deformation, image frames exhibit varying degrees of distortion across different regions. Particularly, objects near the image borders or poles are more susceptible to distortion compared to those in the central regions. This distortion manifests itself as the stretching or compression of objects, leading to inconsistencies in object size and shape throughout the video sequence. Consequently, accurately estimating motion and generating intermediate frames that maintain visual coherence become formidable tasks that require innovative solutions to mitigate the effects of distortion and ensure the fidelity of interpolated frames. Regarding temporal aspects, accurate estimation of temporal motion patterns is also a significant challenge in Omnidirectional Video Interpolation. This includes accurately estimating parameters such as object trajectories, velocities, and accelerations over time within the video sequence. Furthermore, the distortion present in ERP images arises mainly from latitude-dependent deformation, which manifests itself as distortion in the vertical dimension of the image. It’s noteworthy that objects transitioning from regions near the image center (low latitude) to the upper and lower poles (high latitude) undergo alterations in appearance, including variations in area or volume. This deformation phenomenon, known as depth distortion or spherical distortion, is attributed to the distinct representation of spherical images in ERP projection compared to flat images.

### v3

@@ Abstract
Head-mounted 360° displays and portable 360° cameras have significantly progressed, providing viewers a realistic and immersive experience. However, many omnidirectional videos have low frame rates that can lead to visual fatigue, and the prevailing plane frame interpolation methodologies are unsuitable for omnidirectional video interpolation because they are designed solely for traditional videos. This paper introduces the benchmark dataset, 360VFI, for Omnidirectional Video Frame Interpolation. We present a practical implementation that introduces a distortion prior from omnidirectional video into the network to modulate distortions. Specifically, we propose a pyramid distortion-sensitive feature extractor that uses the unique characteristics of equirectangular projection (ERP) format as prior information. Moreover, we devise a decoder that uses an affine transformation to further facilitate the synthesis of intermediate frames. 360VFI is the first dataset and benchmark that explores the challenge of Omnidirectional Video Frame Interpolation. Through our benchmark analysis, we present four different distortion condition scenes in the proposed 360VFI dataset to evaluate the challenges triggered by distortion during interpolation. Besides, experimental results demonstrate that Omnidirectional Video Interpolation can be effectively improved by modeling for omnidirectional distortion.
@@ 1 Introduction
Figure 1: Left: Traditional Video Frame Interpolation. The inputs are two adjacent plane frames, and the output is a target plane frame. Right: Omnidirectional Video Frame Interpolation. The inputs are two adjacent omnidirectional frames with a full field-of-view from an omnidirectional video, and the output is a target omnidirectional frame. Original omnidirectional frames are spherical, and the most common format is the equirectangular projection type (ERP). The two kinds of omnidirectional video formats can be projected from each other, and our proposed method tackles ERP videos.
Traditional Plane Video Frame Interpolation (Plane VFI) techniques have been widely used to address the challenge of low frame rates. Significant progress in these methods has been driven by the development of optical flow networks (Dosovitskiy et al., 2015; Sun et al., 2018; Teed & Deng, 2020; Kong et al., 2021), which allow for accurate frame registration in video sequences by establishing explicit correspondences between frames (Jiang et al., 2018a; Xu et al., 2019; Niklaus & Liu, 2020; Park et al., 2021). These flow-based methods generally follow a three-step process: first, estimating optical flow between two input frames and the target frame; second, warping the input frames or their features using predicted flow fields to achieve spatial alignment; and third, refining these warped frames or features through a synthesis network to generate the final target frame.
Recently, IFRNet (Kong et al., 2022) has advanced this approach by allowing intermediate flows and features to enhance each other, resulting in sharper moving objects and better texture detail. While effective for traditional videos, applying optical flow to omnidirectional video interpolation (ODI) is less efficient. Recent efforts, such as SLOF (Bhandari et al., 2022) and MPF (Li et al., 2022), have adapted optical flow to the omnidirectional domain. SLOF re-projected ERP (equirectangular projection) frames onto a sphere for rotational augmentation, while MPF transformed ERP frames into multiple projection formats and fused them for more precise flow estimation. However, both approaches neglect ERP distortion as prior knowledge, limiting their efficiency in interpolating omnidirectional frames. Since omnidirectional videos (ODVs) are typically stored and transmitted in the ERP format, introducing significant distortions, traditional frame interpolation methods struggle to handle them effectively. Therefore, directly incorporating ERP distortion as prior knowledge in the interpolation process presents a more effective and efficient solution.
@@
Datasets for plane VFI tasks have been sufficiently researched, including Vimeo90K (Xue et al., 2019), UCF101 (Soomro et al., 2012), and SNU-FILM (Choi et al., 2020). Vimeo90K is the most popular dataset for plane VFI, composed of triplet samples. However, Vimeo90K is not designed for motion extent, and this is very important in omnidirectional video interpolation because of its latitude-dependent distortion. SNU-FILM was proposed for the motion-based interpolation benchmark. Both are focused on Plane VFI. Recently, omnidirectional video datasets have caught the attention of the computer community. Several datasets were proposed for omnidirectional video super-resolution, including ODV360 (Wang et al., 2023), 360VDS, and 360UHD (Baniya et al., 2023). They are collected from either YouTube or by themselves. Therefore, there are no omnidirectional video frame Interpolation (Omni-VFI) datasets that have both sufficient and various samples.
@@
To alleviate these problems, we introduce a dataset for omnidirectional video interpolation called 360VFI dataset. We also propose a benchmark (360VFI), which injects the omnidirectional distortion into the network based on the 360VFI dataset. Specifically, in the 360VFI dataset, our evaluation benchmark has four different settings. We classified the triplets based on vertical motion extent due to the latitude-dependent distortion in the omnidirectional frame. In 360VFI Network, we propose the DistortionGuard module to extract less distorted pyramid features from input frames. Moreover, we propose OmniFTB to refine the distorted ERP target frame gradually. Experimental results demonstrate that omnidirectional video interpolation can be effectively improved by modeling for omnidirectional distortion.
Our contribution can be summarized as follows:
  • We present the first omnidirectional video frame interpolation datasets 360VFI. The dataset covers a wide range of content and motion patterns, providing a comprehensive benchmark for evaluating interpolation methods in the omnidirectional domain.
  • We propose the first practical implementation of using distortion priors for omnidirectional video frame interpolation, offering a more accurate and efficient approach to handling the inherent challenges of omnidirectional video processing.
@@ 2.1 Omnidirectional Data Processing
Omnidirectional images and videos have been attracting increasing attention from computer vision and graphics researchers. Due to the special representation of omnidirectional images, dedicated methods were developed for analysis and processing tasks, such as depth estimation (Wang et al., 2020a), salient object detection (Li et al., 2019b; Wu et al., 2022), Visual Localization (Huang et al., 2024), and video stabilization (Tang et al., 2019). With the recent development of immersive technologies, researchers have been working on the specific problems of omnidirectional video-based applications, such as automatic 360° navigation (Kang & Cho, 2019; Hu et al., 2017), omnidirectional video assessment (Li et al., 2019a), and immersive video editing (Nguyen et al., 2017). However, the absence of consistent temporal correspondences poses challenges for various applications, including Omni-VFI. In response to this challenge, our method aims to address the lack of reliable temporal correspondences in omnidirectional video processing. Furthermore, significant research (Esteves et al., 2018; Cohen et al., 2018) has focused on omnidirectional or 360° data, including spherical and ERP data, aimed at learning sphere representations.
@@ 2.2 Video Frame Interpolation
Video frame interpolation aims at generating non-existent frames between two adjacent input frames. In Long et al. (2016), a pioneering learning-based method was proposed that can directly synthesize intermediate frames between two frames. Consequently, frame interpolation methods based on spatially adaptive convolution kernels (Niklaus et al., 2017; Lee et al., 2020; Cheng & Chen, 2021) or pixel phases (Meyer et al., 2015; 2018) were proposed. However, the former leads to large parameters and high complexity, especially when dealing with complex motion. The latter has difficulty handling complex motion due to an inaccurate estimation of phase and amplitude values. After that, many flow-based methods (Jiang et al., 2018b; Xue et al., 2019) use optical flow to guide warping, but the inaccuracy of the predicted optical flow usually causes distortion of the result, so extra measures (Bao et al., 2019; 2018) are usually taken to refine the warped frame. Some methods generate intermediate frames directly, like CAIN (Choi et al., 2020) using channel attention and RIFE (Huang et al., 2022) using distillation to get a good result from one simple architecture. IFRNet (Kong et al., 2022) proposes a novel single encoder-decoder-based method to jointly perform intermediate flow estimation and intermediate feature refinement for efficient VFI, which also performs real-time inference with excellent accuracy. All video frame interpolation methods mentioned above cannot generate well on omnidirectional frames because they are not designed for them. Additionally, to our best knowledge, few methods have been proposed for Omni-VFI.
@@ 2.3 Conditional Modeling and Integration
In the OSRT (Yu et al., 2023), a distortion-aware transformer is introduced for omnidirectional image super-resolution. This approach incorporates ERP distortion priors into the attention mechanism to adjust for distortion. Specifically, OSRT applies deformations to the feature maps by using offsets derived from latitude-dependent conditions, allowing continuous and adaptive modulation. This process enables the network to model and correct ERP distortion effectively. The Spatial Feature Transform (SFT) (Wang et al., 2018) leverages semantic segmentation probability maps as categorical priors to provide valuable conditioning information for image super-resolution. First, the low-resolution input is processed through a semantic segmentation network to generate probability maps, indicating the likelihood of each pixel belonging to different semantic categories (e.g., sky, buildings, vegetation, etc.). Then, the SFT layer uses these probability maps to learn scaling parameters specific to each category, applying them to modulate the intermediate feature maps. This approach enhances the realism and textural detail of the resulting high-resolution image by aligning it with its semantic context.
@@ 2.4 Omnidirectional Datasets
While datasets for plane VFI tasks have been extensively researched, including Vimeo90K (Xue et al., 2019), UCF101 (Soomro et al., 2012), and SNU-FILM (Choi et al., 2020), there is a lack of dedicated datasets for Omni-VFI. Leveraging existing datasets can streamline the benchmark creation process for Omni-VFI, ensuring standardization and reducing the learning curve for researchers. However, no dataset is used directly in the computer vision community for Omni-VFI. Thanks to the contributions of Wang et al. (2023) and S3PO (Baniya et al., 2023), high-resolution omnidirectional video super-resolution datasets ODV360, 360VDS, and 360UHD have been introduced, which can be utilized in Omni-VFI.
@@ 3.2 Proposed 360VFI Dataset
Our 360VFI dataset adopts a similar format as Vimeo90K (Xue et al., 2019). Each sample in our dataset consists of a triplet of video frames. Notably, the first and third frames are designated as input frames for the Omni-VFI model, while the second serves as the ground truth target frame. In the ODV360 (Wang et al., 2023) dataset, each video comprises 100 frames, while in the 360VDS and 360UHD (Baniya et al., 2023) dataset, the videos consist of 20 frames. We omit the final frame from each video in the ODV360 dataset, resulting in 99 frames per video, subsequently divided into 33 triplets. For videos in the 360VDS and 360UHD datasets, the first and last frames are discarded, leaving 18 frames per video, which are divided into six triplets. Each triplet constitutes an independent sample and the whole dataset consists of 930 triplets. This comprehensive arrangement enables a holistic evaluation of interpolation algorithms against genuine ODV content. We randomly divide these videos into training and testing sets.
@@
Videos in our dataset cover various scenarios, including natural landscapes, playgrounds, interiors of houses and cars, and indoor markets, as shown in Figure 2. To facilitate nuanced evaluations and benchmarking, we stratified the dataset into four distinct settings based on the varying degrees of motion inherent within the omnidirectional scenes, ranging from 0.19 to 9.69. These settings are categorized as easy, middle, hard, and extreme; the triplet numbers and their respective storage are illustrated in Table 2. The motion ranges for these settings are as follows: [0.19, 2], [2, 3], [3, 4], and [4, 9.69]. The frames and optical flow of different partitions are illustrated in Figure 3. This stratification allows researchers to systematically assess the performance of their models across varying levels of motion complexity, facilitating a deeper understanding of model robustness and generalization capabilities. In Table $1$VFI: Video Frame Interpolation; VSR: Video Super-Resolution; OLSE: Outdoor Light Source Estimation; OD: Object Detection, we compare the 360VFI dataset with other datasets. These datasets include plane video, 360° video, and 360° images. Our dataset is the first one composed of 360° video for omnidirectional video interpolation. Our dataset establishes a solid foundation for future research in Omni-VFI. By amalgamating diverse datasets and incorporating motion-based stratification, it not only enriches the existing repository of ODV data but also sets a precedent for structured evaluation and benchmarking in this evolving field. We anticipate that our dataset will serve as a catalyst for innovation and foster the development of more sophisticated and resilient omnidirectional video interpolation models in the future.
@@ 4 Method
where $\Psi$ defines the prior upon which the mapping function can condition. As mentioned in Section 2.3, there are many works using prior knowledge to guide their task in other computer vision tasks. Hence, we propose 360VFI Net, including DistortionGuard and OmniFTB.
4.1 Distortions in Omnidirectional Video
@@
ERP (Equirectangular Projection) is a commonly used method for mapping spherical surfaces onto a plane. In ERP, every point on the sphere is projected in such a way that each line from the polar center maintains the same distance. As a result, the distortion in ERP is latitude-dependent, with all pixels at the same height experiencing the same level of distortion. This makes ERP a practical choice for storing or transmitting omnidirectional images and videos, as it preserves depth and distance perception while minimizing distortions. However, the degree of distortion varies depending on the projection type, and ERP introduces significant distortions, especially at higher latitudes.
To quantify the extent of distortion in ERP, we follow the definition of the stretching ratio ($\mathbf{K}$) as described in Sun et al. (2017a), which measures the distortion at various points in relation to the ideal spherical surface. The stretching ratio $\mathbf{K}$ is determined by the variation in area between the spherical surface and the projection plane. In ERP, the coordinates are defined as $x$ and $y$, and the stretching ratio can be expressed as:
@@
where $\delta S$ represents the area on the spherical surface and $\delta P$ represents the area on the projection plane, with $x\in(-\pi,\pi)$ and $y\in(-\frac{\pi}{2},\frac{\pi}{2})$. From this equation, it becomes clear that ERP distortion depends solely on latitude, with the most severe distortion occurring in the polar regions.
For an input frame $X\in\mathbb{R}^{C\times M\times N}$, we can derive the distortion condition map $C_{d}\in\mathbb{R}^{1\times M\times N}$ as follows:
@@
where $m$ is the height index of the input frame. This map helps capture the latitude-dependent distortion across the frame.
Spatial distortion is a major challenge in omnidirectional video interpolation due to the projection methods used, such as ERP. Objects near the image borders or poles suffer from significant stretching or compression, leading to inconsistencies in their size and shape across the video sequence. This latitude-dependent deformation complicates both the motion estimation and frame interpolation processes, making it difficult to maintain visual coherence. Temporal motion patterns also present a challenge in omnidirectional video interpolation. Accurately estimating object trajectories, velocities, and accelerations within the video sequence is complicated by the distortion present in ERP images, particularly as objects move from the center (low latitude) to the poles (high latitude). This distortion, referred to as depth distortion or spherical distortion, alters the appearance and size of objects, further complicating interpolation efforts.